Iron produces from fe2o3 by reaction with carbon. The reaction products are




Fe, FeO and CO. If we mix 100 Kg of Fe2o3 and 24 Kg of C and produced 12 Kg of FeO, determine:




a) Limiting reaction and excess reaction.




b) % Excess reaction.




c) % Conversion of Fe2o3 and C.




d) The degree of completion of reaction.




e) Mass of Fe in Kg.




Fe2o3+ 2C / Fe+FeO+2CO




Mw;




Fe = 56




C=12.




O = 16

1
Expert's answer
2022-03-08T14:54:03-0500

Fe2O3 + 2C → Fe + FeO + 2CO

a)      Limiting reagent:

Molar mass Fe2O3: 159.69

Moles Fe2O3: 100 / 159.69 = 0.626 kmol

Molar mass C: 12.01

Moles Fe2O3: 24 / 12.01 = 2.00 kmol

The mole ratio according to the equation: 1 (Fe2O3) to 2 (C); meaning 2.00 mole C would require 1.00 mole (Fe2O3). As we have less (0.626), we conclude that Fe2O3 is limiting reagent.

  

b)     % Excess reagent:

C is in excess; thus, it is an excess reagent.

 

c)      % Conversion of Fe2O3 and C:

100% of Fe2O3 is used up, because it is a limiting reagent;

0.626 kmol x 2 = 1,252 kmol (C) reacts

1.252 / 2.00 = 0,626 or 62,6% - converted C.

 

d)     The degree of completion of reaction:

Molar mass FeO: 71.84

Moles FeO: 12 / 71.84 = 0.167 kmol

Theoretically we can obtain: 0.626 kmol (same as Fe2O3)

0.167 / 0.626 = 0.27 or 27% - the degree of completion.

 

e)     Mass of Fe in kg:

Moles of Fe: 0.167 kmol

Molar mass Fe: 55.85

Mass Fe: 55.85 x 0.167 = 9.33 kg.